Suppose there are only two goods in the economy, apples and oranges and the market that composes CPI is 4 apples 2 oranges. Assume that 2011 is the base year.

2011 2012

P apple 1 1.03

P Orange 1 1.09

Question?

Consider an individual who receives a fixed income from the government. Suppose this person receives 24$ and uses it to buy 16 apples and 8 oranges. Suppose their income is indexed to the CPI. How much will they receive on 2012? Since oranges are more costly than apples on January 2012, the fixed income might decide to consume more apples? Suppose they chose to consume 18 apples. How many oranges will they consume?

PLease and thank you :)
Answer #1

Ans

Sum of old price=1+1=2

Sum of new price=1.03+1.09=2.12

Cpi=2.12/2(100)=106

So prices have risen by 106-100=6%

Hence income should increase by =24+24(6%)=25.44. Thus is amount that they will receive in 2012

Amount left after consuming 18 apples=25.44-(1.03)(18)=6.9

So they will consume oranges=6.9/1.09=6.33